a 7.0 mm diameter copper wire carries a current of 0.7ua.what is the current density?

1
Expert's answer
2022-03-28T14:25:45-0400

"I=0.7\mu A\\r=3.5mm"

"J=\frac{I}{A}=\frac{I}{\pi r^2}"

"J=\frac{0.7\times10^{-6}}{3.14\times(3.5\times10^{-3})^2}"

"J=18.19\times10^{-3}A/m^2"